Use two unit multipliers to convert 4.4 square meters to square centimeters.
Ive tried to do it several ways and I cannot figure out how to go from meters squared to centimeters squared?

Answer by oberobic(2304)

1 m = 100 cm
so
1 m^2 = 100 cm * 100 cm = 10000 cm^2
.
4.4 m^2 = 4.4 m^2 *10000 cm^2/m^2 = 44000 cm^2